OpenStudy (anonymous):

f(x)=x+3 g(x)=x-3 find fog(-3)

OpenStudy (anonymous):

f(x)=x+3 g(x)=x-3 f(g(x)) = put x in f(x) = g(x) f(g(x)) = (x-3) +3 f(g(x)) = x fog(-3) = ?

OpenStudy (anonymous):

I keep coming up with -3, but the answer is supposed to be 9

OpenStudy (anonymous):

g(x)=-3-3=-6 f(x)= -6+3=-3

OpenStudy (anonymous):

it should be -3. I dont know any other way http://www.purplemath.com/modules/fcncomp2.htm

OpenStudy (anonymous):

Ok thanks! I thought I was going crazy.

OpenStudy (anonymous):

anytime :)